To buy both a new car and a new house, Tina sought two loans totalling $44,980. The simple interest rate on the first loan was 1.8%, while the simple interest rate on the second loan was 1.2%. At the end of the first year, Tina paid a combined interest payment of $679.64. What were the amounts of the two loans?
Answer by stanbon(75887)

To buy both a new car and a new house, Tina sought two loans totaling $44,980. The simple interest rate on the first loan was 1.8%, while the simple interest rate on the second loan was 1.2%. At the end of the first year, Tina paid a combined interest payment of $679.64. What were the amounts of the two loans?
--------------------
Equations:
Quantity :: x + y = 44900
Interest :0.018x+0.012y = 679.64
-------------------------------------
Multiply thru the quantity Eq. by 18
Multiply thru the interest Eq. by 1000
------
18x + 18y = 18*44900
18x + 12y = 679640
------------------------
Subtract and solve for "y":
6y = 128560
y = 21426.67 (amt. of 1.2% loan)
---
Solve for "x":
x + y = 44900
x + 21426.67 = 44900
x = 23473.33 (amt. of 1.8% loan)
===================================
Cheers,
Stan H.
===========